# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v4

@@ 4.4 Results
Figure 3: Simulation results for the nonlinear and binary continuous outcomes with sample sizes (a) $N=40$ and (b) $N=200$. In the Bias boxplots, the black dots represent the medians. In the CPplots, the red dashed line indicates the 95% coverage level, while the gray dashed lines indicate 93% and 97% coverage levels. In the Powerplots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power. Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
@@
Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels under each outcome. We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
@@
From Figure 3, we observe that when the sample size is very small ($N=40$), there are many outliers in the boxplots for the “No selection” case for all covariate-adjusted estimators. This is expected, since they include a relatively large number of covariates compared to the sample size. Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators exhibit wider bias box widths, with the Simple estimator widest. The Corr method does not exhibit consistent performance across different parameter choices of $k$ and $\xi$. For example, Corr with $k=1$ or $k=3$ is notably less efficient when the sample size is $N=200$, yet performs relatively well under $N=40$. In contrast, Corr with $\xi=0.10$ shows poor performance under $N=40$ but improves substantially with a larger sample size of $N=200$. Lasso and adaptive Lasso exhibit similar performance patterns under the same estimator and are overall the most efficient and consistently performing methods across simulation settings. An exception is observed with the ANHECOVA estimator under $N=40$ for both outcomes, where some outliers appear. Nevertheless, in all other scenarios—particularly when paired with the AIPW estimator—these methods consistently demonstrate stable gains in efficiency and statistical power. These findings suggest that Simple, Pre-test, and Corr methods are generally less efficient in our simulation setting, particularly when compared to Lasso and adaptive Lasso.
@@
In terms of CPand power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size. However, even when $N=200$, we observe below nominal or missing CPfor all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CPunder the continuous outcome with $N=200$. However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material. This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research. Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
@@
Futhermore, although we do not include the results of other DGPs, we briefly summarize them here. The results for other two outcomes (linear continuous and linear binary) are similar to those shown in Figure 3 and our comments above. At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CPvalues that are closer to the 95% nominal level across all outcomes. However, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA without variable selection persists even at $N=500$. This indicates that, despite a seemingly adequate sample size relative to the data dimension, the sandwich variance estimator may remain computationally unstable or fail to converge when the number of covariates is large. This finding further highlights the advantage of applying variable selection to control the effective dimensionality, thereby facilitating the computation of model-based variance estimators.
@@ 5 Case Study
We apply the COADVISE framework to analyze the Best Apnea Interventions for Research (BestAIR) trial data, available through the National Sleep Research Resources (NSRR) (https://sleepdata.org/datasets/bestair/files/datasets). The dataset includes 169 participants, with 83 assigned to the active treatment group (continuous positive airway pressure [CPAP]) and 86 to a combined control arm. The primary objective of the study is to assess the impact of CPAP on 24-hour systolic blood pressure (SBP). The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$). We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high (130$\geq$ 130≥ 130 mmHg) or low (<130<130< 130 mmHg) based on a 130 mmHg threshold.
@@ 6.2 Discussion
Furthermore, our framework can be extended to a wide range of more general and complex settings. These include covariate-adaptive randomization (Bannick et al., 2023; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020). Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025). Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2023; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
@@ Acknowledgments
S.1 Summary of data imputation methods in COADVISE
Imputation method
Handles missing
covariates
Handles missing
outcome
Assumption
@@
Multiple imputation by chained equations
(mice; Van Buuren and Groothuis-Oudshoorn (2011))
$\checkmark$ $\checkmark$ MAR
Random forest
(missForest; Stekhoven and Bühlmann (2012))
$\checkmark$ $\checkmark$ MAR
Inverse probability weighting
(ipw; Robins et al. (1994))
$\times$ $\checkmark$ MAR
Missingness indicator
(missInd; Zhao et al. (2024);
Zhao and Ding (2022); Song et al. (2021))
$\checkmark$ $\times$ MAR
@@ X The covariate vector/matrix/data frame $\checkmark$
[…]
@@
#### We first use Coadvise to perform a variable selection, using the Corr.k method with k=5, and results are saved in the following "step1" object
@@
#### We then use covariates selected in step1 for ANCOVA for further analyses
@@
rc.simple <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
@@
0 0 $=\frac{1-\pi_{1}}{\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{1}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})[Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})]\right\},$
0 0 $=\frac{\pi_{1}}{1-\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{0}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})[Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})]\right\},$ (B.1.4)
@@
[…]
@@
$\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}Y(1)^{2}+\frac{\pi_{1}}{1-\pi_{1}}Y(0)^{2}\right\}-\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}$
@@
For both $a=0,1$, we note that $I_{a}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\{I_{a}(\mathbf{X})\}$, where $I_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]$. Therefore, $I_{a}(\mathbf{X})$ achieves its maximum when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})=\mathbb{E}(Y(a)\mid\mathbf{X})$. This implies that when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ is the fully specified model for $\mathbb{E}(Y(a)\mid\mathbf{X})$, $\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ reaches its maximal value $\mathbb{E}\{\mathbb{E}\{Y(a)^{2}\mid\mathbf{X}\}\}=\mathbb{E}\{Y(a)^{2}\}$, corresponding to the highest efficiency gain. Additionally, in this case, the second term in the (S.1) of the squared loss is always zero, which ensures $\boldsymbol{\beta}_{a}^{*}$ is the variance minimizer. However, this is typically unrealistic in practice, and if the model is not well-specified, it is possible that (S.3) can be negative. In the following, we outline some conditions for achieving efficiency gains.
@@
[…]
Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality). Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from [0.11,)[0.11$\infty$)[ 0.11 , ∞ ). When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
where $\boldsymbol{I}_{50}$ is the $\times 50$ 50 identity matrix.
@@
In Figures S.2.1S.2.16, we present full results (Bias, CPand Power%) of simulation in Section 4 of the main text.
@@
Figure S.2.1: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=40$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.2: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.3: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=200$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.4: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.5: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.6: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.7: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.8: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.9: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=40$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.

### v5

@@ 4.4 Results
Figure 3: Simulation results for nonlinear and binary outcomes with sample sizes (a) $N=40$ and (b) $N=200$. In the Bias boxplots, black dots indicate medians. In the CP% plots, the red dashed lines mark the 95% coverage level, while the gray dashed lines mark the 93% and 97% levels. In the Power% plots, the red dashed lines indicate 100% power, and the gray dashed lines indicate 80% power. Missing values in CP% and Power% for ANHECOVA and ANCOVA arise from the non-obtainable sandwich variance estimators in high-dimensional settings.
@@
Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CP% and Power% in the lower four panels under each outcome. We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
@@
From Figure 3, we observe that when the sample size is very small ($N=40$), there are many outliers in the boxplots for the “No selection” case for all covariate-adjusted estimators. This is expected, since they include a relatively large number of covariates compared to the sample size. Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and $0.1$) estimators exhibit wider bias box widths, with the Simple estimator widest. The Corr method does not exhibit consistent performance across different parameter choices of $k$ and $\xi$. For example, Corr with $k=1$ or $k=3$ is notably less efficient when the sample size is $N=200$, yet performs relatively well under $N=40$. In contrast, Corr with $\xi=0.10$ shows poor performance under $N=40$ but improves substantially with a larger sample size of $N=200$. Lasso and adaptive Lasso exhibit similar performance patterns under the same estimator and are overall the most efficient and consistently performing methods across simulation settings. An exception is observed with the ANHECOVA estimator under $N=40$ for both outcomes, where some outliers appear. Nevertheless, in all other scenarios—particularly when paired with the AIPW estimator—these methods consistently demonstrate stable gains in efficiency and statistical power. These findings suggest that Simple, Pre-test, and Corr methods are generally less efficient in our simulation setting, particularly when compared to Lasso and adaptive Lasso.
@@
In terms of CP% and power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size. However, even when $N=200$, we observe below nominal or missing CP% for all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CP% under the continuous outcome with $N=200$. However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material. This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research. Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
@@
Futhermore, although we do not include the results of other DGPs, we briefly summarize them here. The results for other two outcomes (linear continuous and linear binary) are similar to those shown in Figure 3 and our comments above. At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CP% values that are closer to the 95% nominal level across all outcomes. However, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA without variable selection persists even at $N=500$. This indicates that, despite a seemingly adequate sample size relative to the data dimension, the sandwich variance estimator may remain computationally unstable or fail to converge when the number of covariates is large. This finding further highlights the advantage of applying variable selection to control the effective dimensionality, thereby facilitating the computation of model-based variance estimators.
@@ 5 Case Study
We apply the COADVISE framework to analyze the Best Apnea Interventions for Research (BestAIR) trial data, available through the National Sleep Research Resources (NSRR) (https://sleepdata.org/datasets/bestair/files/datasets). The dataset includes 169 participants, with 83 assigned to the active treatment group (continuous positive airway pressure [CPAP]) and 86 to a combined control arm. The primary objective of the study is to assess the impact of CPAP on 24-hour systolic blood pressure (SBP). The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$). We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high ($\geq 130$ mmHg) or low (<130<130 mmHg) based on a 130 mmHg threshold.
@@ 6.2 Discussion
Furthermore, our framework can be extended to a wide range of more general and complex settings. These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020). Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025a). Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
@@ Acknowledgments
S.1 Summary of missing data imputation methods in COADVISE
Imputation method Handles missing covariates Handles missing outcome Assumption
@@
Multiple imputation by chained equations (mice; Van Buuren and Groothuis-Oudshoorn (2011)) $\checkmark$ $\checkmark$ MAR
Random forest (missForest; Stekhoven and Bühlmann (2012)) $\checkmark$ $\checkmark$ MAR
Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
@@ X The covariate vector/matrix/data frame $\checkmark$
lin.cal Whether to use linear calibration in AIPW $\times$ FALSE TRUE, FALSE
@@
Let $\mathcal{R}^{p}$ denote the $p$-dimensional Euclidean space $(p>0)$. Define $N_{a}=\sum_{i=1}^{N}I(A_{i}=a)$ as the number of units with treatment assignment $A_{i}=a$, $s_{N}^{2}(\mathcal{W})$ as the sample variance of the i.i.d. data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$, and $\text{cov}_{N}(\mathcal{W},\mathcal{V})$ as the sample covariance of the data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{N}\}$. Similarly, $s_{N_{a}}^{2}(\mathcal{W})$ represents the sample variance for all $W_{i}\in\mathcal{W}$ with $A_{i}=a$, and $\text{cov}_{N_{a}}(\mathcal{W},\mathcal{V})$ denotes the sample covariance for all $W_{i}\in\mathcal{W}$ and $V_{i}\in\mathcal{V}$ with $A_{i}=a$, where $a=0,1$. Denote $\mathbb{P}_{N}[f(W)]=N^{-1}\sum_{i=1}^{N}f(W_{i})$ the empirical mean of any function of the i.i.d. observed data $\mathcal{W}$.
@@
#### We first use Coadvise to perform variable selection, using the Corr.k method with k=5, and results are saved in the "step1" object
@@
#### We then use covariates selected in step1 for ANCOVA
@@
rc.simple <- robincar_linear(df=df.RobinCar, treat_col="A",
response_col="Y",
@@
$0$ $=\frac{1-\pi_{1}}{\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{1}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})[Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})]\right\},$
$0$ $=\frac{\pi_{1}}{1-\pi_{1}}\mathbb{E}\left\{g_{\boldsymbol{\beta}_{0}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})[Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})]\right\},$ (B.1.4)
@@
Furthermore, this result can easily be extended situations where a variable selection step is performed before AIPW estimation. Suppose we select $s$ variables from $\mathbf{X}$ for $Y(a)$ model, where $0<s<p$. When fitting the OLS estimator, the only difference is that $\boldsymbol{\beta}\in\mathcal{R}^{s}$ instead of $\mathcal{R}^{p}$ for estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Here, $\widehat{\tau}_{\text{simple}}$ remains a special case of estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Therefore, the proof remains straightforward.
@@
$\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ $=\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}Y(1)^{2}+\frac{\pi_{1}}{1-\pi_{1}}Y(0)^{2}\right\}$
$\qquad-\mathbb{E}\left\{\frac{1-\pi_{1}}{\pi_{1}}\{Y(1)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{1}^{*})\}^{2}+\frac{\pi_{1}}{1-\pi_{1}}\{Y(0)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}^{*})\}^{2}\right\}$
@@
For $a=0,1$, we note that $I_{a}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2Y(a)-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\left\{g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]\right\}=\mathbb{E}\{I_{a}(\mathbf{X})\}$, where $I_{a}(\mathbf{X})=g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})[2\mathbb{E}(Y(a)\mid\mathbf{X})-g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})]$. Therefore, $I_{a}(\mathbf{X})$ achieves its maximum when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})=\mathbb{E}(Y(a)\mid\mathbf{X})$. This implies that when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})$ is the fully specified model for $\mathbb{E}(Y(a)\mid\mathbf{X})$, $\mathbb{V}(\widehat{\tau}_{\text{simple}})-\mathbb{V}(\widehat{\tau}^{g})$ reaches its maximal value $\mathbb{E}\{\mathbb{E}\{Y(a)^{2}\mid\mathbf{X}\}\}=\mathbb{E}\{Y(a)^{2}\}$, corresponding to the highest efficiency gain. Additionally, in this case, the second term in the (S.1) of the squared loss is always zero, which ensures $\boldsymbol{\beta}_{a}^{*}$ is the variance minimizer. However, this is typically unrealistic in practice, and if the model is not well-specified, it is possible that (S.3) can be negative. In the following, we outline some conditions for achieving efficiency gains.
@@
Figure S.3.1 illustrates the right-hand side of inequality in (ii). From the figure, if $\boldsymbol{\beta}_{a}^{*}$ is such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 1/2$, the (absolute value of) bias is allowed to be relatively large. However, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ is either smaller or larger, the bias must be smaller, approaching zero when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 0$ or $1$. In other words, when the true proportion of $Y(a)$ is very small or vary large, the model must better approximate the true value. But when the true proportion is around $50\%$, more error in the posited model is acceptable. The intuition is when the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ approaches a medium value, the uncertainty regarding whether the true $Y(a)$ is 0 or $1$ increases.
Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality). Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from $0.11,\infty)$. When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
where $\boldsymbol{I}_{50}$ is the $50\times 50$ identity matrix.
@@
In Figures S.2.1S.2.16, we present full results (Bias, CP% and Power%) of simulation in Section 4 of the main text.
@@
Figure S.2.1: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=40$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.2: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.3: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=200$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.4: Simulation results under continuous outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.5: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.6: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.7: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.8: Simulation results under continuous outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.9: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=40$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.